OpenStudy (alystar):

A bin contains 6 purple pens, 5 red pens, and 4 green pens. You randomly select a pen. What is the probability of selecting a red pen? the answer to that is 1/3 Using the fully-stocked bin from the above question, what is the probability of selecting a green pen on Monday, then selecting a purple one on Wednesday, to replace the green one you lost Tuesday? a) 3/7 b) 4/35 c) 8/75 d) 3/4

OpenStudy (alystar):

I just need a refresher on how to do this :)

OpenStudy (cruffo):

On Monday: fully stocked: 6 purple, 5 red, 4 green = total of 15 pens. Prob that you choose a green pen: 4 out of 15 = 4/15 On Wednesday: not fully stocked.. 6 purple, 5 red, 3 green = total of 14 pens Prob that you choose a purple pen: 6 out of 14 = 6/14 = 3/7 And probabilities multiply : probability of green pen then purple (without replacement of original green pen) = (4/15) * (3/7) = (4 * 3) / (15 * 7) = 4/35

OpenStudy (alystar):

Thanks @cruffo

OpenStudy (cruffo):

Cheers!